A Bibliography of Publications of Igor Jurisica

Igor Jurisica University of Toronto

Toronto, Ontario M5S 1A4

Canada Tel: +1 416 978 7589 FAX: +1 416 978 1455

E-mail: juris@cs.utoronto.ca (Internet)

25 November 2011 Version 1.04

Abstract

This bibliography records publications of Igor Jurisica.

Title word cross-reference

12-16 [AAA92]. AAAI [AAA92]. AAAI-92 [AAA92]. Applied [Jur94a]. approach [GJ92, Jur92]. Artificial [AAA92]. assessment [Jur94e]. base [Jur92]. Based [Jur94a, Jur94d, Jur93]. CA [AAA92]. Case [Jur94d, Jur93]. Case-Based [Jur94d, Jur93]. Cases [Jur94a]. Conference [AAA92]. Context [Jur94a]. Context-Based [Jur94a]. EBL [GJ92]. Features [Jur94b]. information [Jur94c]. Intelligence [AAA92]. Issues [Jur94d, Jur93]. July [AAA92]. knowledge [Jur92]. learning [Jur92]. machine [Jur92]. Management [Jur94d, Jur92, Jur93]. Menlo [AAA92]. National [AAA92]. optimization [Jur92]. Park [AAA92]. problem [GJ92]. Proceedings [AAA92]. Query [Jur92]. Reasoning [Jur94d, Jur93]. Relevant [Jur94a, Jur94c]. Representation [Jur93, Jur94d]. Retrieval [Jur94a]. retrieve [Jur94c].

Similarity [Jur94a, Jur94e, Jur94b]. solving [GJ92]. statistical [GJ92]. Systems [Jur94d, Jur92].

Tenth [AAA92].

utility [GJ92].

References

AAAI:1992:PAT

[AAA92] Proceedings of AAAI-92: Tenth National Conference on Artificial Intelligence, July 12-16, 1992. Menlo Park, CA. AAAI Press and MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-51063-4. LCCN Q334.N36 1992.

Greiner:1992:SAS

[GJ92] Russell Greiner and Igor Jurišica. A statistical approach to solving the EBL utility problem. In AAAI-92 [AAA92], pages 241–247. ISBN 0-262-51063-4. LCCN Q334.N36 1992.

Jurisica:1992:QOK

[Jur92] Igor Jurišica. Query optimization for knowledge base management systems; A machine learning approach. Master’s thesis, Department of Computer Science, University of Toronto, Toronto, Ontario, 1992.

Jurisica:1993:RMI

[Jur93] Igor Jurišica. Representation and management issues for case-based reasoning. Depth Paper, Department of Computer Science, University of Toronto, Toronto, Canada,

Jurisica:1994:CBS

[Jur94a] Igor Jurišica. Context-based similarity applied to retrieval of relevant cases. Technical Report DKBS-TR-94-5, University of Toronto, Department of Computer Science, Toronto, Ontario, 1994.

Jurisica:1994:FS

[Jur94b] Igor Jurišica. Features of similarity. In 4^{th} Annual IRIS/PRECARN Conference. Toronto, ON, June 21-23 1994.

Jurisica:1994:HRR

[Jur94c] Igor Jurišica. How to retrieve relevant information? In Russell Greiner, editor, Proc. of the AAAI Fall Symposium Series on Relevance, pages 101–104. New Orleans, Louisiana, 1994.

Jurisica:1994:RMI

[Jur94d] Igor Jurišica. Representation and management issues for case-based reasoning systems. In TRIO/ITRC Research Retreat. Queen’s University, Kingston, May 10-12 1994.

Jurisica:1994:SA

[Jur94e] Igor Jurišica. Similarity assessment. In TRIO/ITRC Research Retreat. Queen’s University, Kingston, May 10-12 1994.